ryan and karl spent 28 hours building a tree house. ryan worked 4 more hours than karl. how many hours did each work.
Answer by mananth(16946)

let karl work for x hours
Ryan works for x+4 hours
..
x+x+4 = 28
2x+4=28
add -4
2x+4-4= 28-4
2x=24
/2
x=12 --- time karl worked
Ryan = 12+4 = 16 hours
...
m.ananth@hotmail.ca